INFRARED ACTION SPECTROSCOPY AND STABILITY OF THE OH-HONO_{2} REACTION INTERME-DIATE

\underline{BRIDGET A. O’DONNELL}, EUNICE X. J. LI, and MARSHA I. LESTER, Department of Chemistry, University of Pennsylvania, Philadelphia, PA 19104-6323.

The reaction of nitric acid (HONO_{2}) with the hydroxyl radical (OH) is of significant atmospheric interest. Kinetic studies have revealed unusual behavior including a negative temperature dependence, pressure dependence, and an overall reaction rate strongly affected by isotopic substitution.^{a} This suggests that the reaction occurs through an intermediate, theoretically predicted to be a hydrogen-bonded OH-HONO_{2} complex in a six-membered ring-like configuration.^{b} In this study, OH-HONO_{2} is produced in the gas phase by the association of photolytically generated OH radicals with HONO_{2} in the collisional region of a supersonic expansion. Infrared action spectroscopy is used to identify the fundamental OH radical stretch (ν_{1}) and OH nitric acid stretch (ν_{2}) of the complex. The rotationally structured ν_{1} band^{c} is centered at 3216 cm^{−1}, while the extensively broadened ν_{2} band is centered at 3260 cm^{−1}, both shifted from their respective monomer. Following infrared excitation, the OH fragments resulting from vibrational predissociation are detected by laser-induced fluorescence on the OH A^{2}Σ^{+}-X^{2}Π (1,0) transition. The highest observed OH (v=0) product channels, J =17/2, Ω=1/2 for the OH radical stretch and J =15/2, Ω=1/2 for the OH nitric acid stretch, lead to determination of an upper limit for the binding energy of OH-HONO_{2}, D_{0}≤6.2 kcal mol^{−1}, which is consistent with ab initio predictions.^{b}

aS. S. Brown, R. K. Talukdar, and A. R. Ravishankara, J. Phys. Chem. A 105, 9 (2001).

bS. Aloisio and J. S. Francisco, J. Phys. Chem. A 103, 30 (1999).

cOpen-shell theoretical analysis by Mark D. Marshall.